### v1

# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

Abstract.
This paper investigates the neural dropout method as a post-processing bias mitigation for deep neural networks (DNNs). Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications. While neural networks are exceptionally good at finding statistical patterns from data, they are notorious for over-fitting the training datasets that may encode and amplify existing biases from the historical data. Existing bias mitigation algorithms often require either modifying the input dataset or modifying the learning algorithms. We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve fairness of pre-trained DNNs. However, finding the ideal set of neurons to drop is a combinatorial problem.
We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference after training. Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility. We show that our design of randomized algorithms are effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation. We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of search algorithms on the results. Finally, we empirically and conceptually compare NeuFair to different state-of-the-art bias mitigators.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austria
1. Introduction
Artificial intelligence (AI), increasingly deployed with deep neural network (DNN) components, has become an integral part of modern software solutions that assist in socio-economic and legal-critical decision-making processes such as releasing patients (Kharpal, 2018), identifying loan defaults (Group, 2018), and detecting tax evasion (Tankersley, 2023).
Despite many advances made possible by AI, there are challenges that require understanding the dimensions and implications of deploying AI-driven software solutions. One such concern about the trustworthiness of AI is discrimination. Unfortunately, there are plenty of fairness defects in real systems. Parole decision-making software was found to harm black and Hispanic defendants by falsely predicting a higher risk of recidivism than for non-Hispanic white defendants (Julia Angwin and Kirchne, 2021); Amazon’s hiring algorithm disproportionately rejected more female applicants than male applicants (Ingold and Soper, 2016); and data-driven auditing algorithm selected black taxpayers with earned income tax credit claims (EITC) at much higher rates than other racial groups for an audit (Tankersley, 2023). As evidenced by these examples, resulting software may particularly disadvantage minorities and protected groups and be found non-compliant with law such as the US Civil Rights Act (Blumrosen, 1978). Hence, helping programmers and users to mitigate unfairness in social-critical data-driven software systems is crucial to ensure inclusion in our modern, increasingly digital society.
The software engineering (SE) community has spent significant efforts to address discrimination in the automated data-driven software solutions (Galhotra et al., 2017; Aggarwal et al., 2019; Udeshi et al., 2018; Zhang et al., 2020a; Chakraborty et al., 2020). Fairness has been treated as a critical meta-property that requires analysis beyond functional correctness and measurements beyond prediction accuracy (Brun and Meliou, 2018). Thus, the community presents various testing (Galhotra et al., 2017; Aggarwal et al., 2019; Zhang et al., 2020b), debugging (Gohar et al., 2023; Monjezi et al., 2023; Yu et al., 2023), and mitigation (Zhang and Harman, 2021; Biswas and Rajan, 2021) techniques to address fairness defects in data-driven software.
Broadly, fairness mitigation can be applied in pre-processing stage (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing step (e.g., changing the loss function to include fairness constraints during training), or post-processing stage (e.g., changing the logic of a pre-trained model). However, when the decision logic of AI is encoded via deep neural networks, it becomes increasingly challenging to mitigate unfairness due to its black-box uninterpretable nature.
Since the DNN-based software development is derived by principle of information bottleneck (Tishby et al., 2000), which provides a trade-off between over-fitting and parsimonious signals; we posit that over-fitted DNN models may contribute and amplify historical biases from the training datasets. Therefore, prevalent techniques such as dropout methods (Srivastava et al., 2014; Srivastava, 2013)—a process of randomly dropping neurons during training—might be effective in mitigating unfairness in the pre-trained DNN models. While the dropout strategy has been significantly used to prevent over-fitting during the training of DNNs, to the best of our knowledge, this is the first work to systematically leverage dropout methods for mitigating unfairness in pre-trained DNN models, as a post-processing mitigation method.
However, finding the optimal subset of dropout neurons is a combinatorial problem that requires an exhaustive search in the subset of all combinations of neuron dropouts, and hence is intractable. To overcome the computational challenges, we explore the class of randomized algorithms with Markov chain Monte Carlo (MCMC) strategies to efficiently explore the search space with statistical guarantees. In doing so, we pose the following research questions:
RQ1. How successful are randomized algorithms in fairness repairing of DNNs via dropouts?
RQ2. Are there dropout strategies that improve model fairness and utility together?
RQ3. What are the design considerations of search algorithms for efficient and effective dropouts?
RQ4. How do dropout strategies compare to the state-of-the-art post-processing (bias) mitigators?
To answer these research questions, we present NeuFair (Neural Network Fairness Repair): a set of randomized search algorithms to improve the fairness of DNNs via inference-time neuron dropouts. We design and implement simulated annealing (SA) and random walk (RW) strategies that efficiently explore the state-space of neuron dropouts where we encode the frontiers of fairness and utility in a cost function.
We evaluate NeuFair over $7$ deep neural network benchmarks as trained over $5$ socially critical applications with significant fairness implications. We found that randomized strategies of neuron drops can successfully improve fairness (up to 69%) with a minimal utility degradation in the vast majority of cases. We also report a pathological case and reasons behind a failure of NeuFair. We also observe that NeuFair can improve fairness and utility together and provide intuitive explanations of such phenomena. Furthermore, we examine different hyperparameter configuration options of randomized algorithms. While some hyperparameters always influence fairness with positive or negative impacts; we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark. Finally, we show the effectiveness of the SA algorithm, compared to the RW and the state-of-the-art post-processing (bias) mitigator (Monjezi et al., 2023). In summary, the key contributions of this paper are:
  1. (1)
    Inference-Time dropout for fairness. To the best of our knowledge, we present the first dropout method of bias mitigation over pre-trained deep neural networks,
  2. (2)
    Randomized algorithms for fairness. We create a well-defined framework to formulate the combinatorial inference-time dropout problem in DNNs using randomized algorithms,
  3. (3)
    Experimental evaluations. We implement the randomized algorithms in NeuFair and evaluate its effectiveness and efficiency vis-a-vis the state-of-the-art techniques.
2. Background
We provide a background on the model utility and fairness measures. Then, we briefly discuss various bias mitigation methods from the AI fairness literature.
2.1. Notions of Model Utility
[…]
$\text{Accuracy}_{h}=\frac{TP+TN}{TP+TN+FP+FN}$
[…]
Accuracy can be used to gauge the overall performance of a classifier. However, for imbalanced datasets used in fairness evaluations, accuracy is a poor metric as the number of negative samples far outweighs the positive samples. For example, predicting all samples as negative (0 true positives) in the Bank (Dua and Graff, 2017a) dataset yields an accuracy of 88% but the F1 score would be 0/undefined. Since F1 is defined as the harmonic mean of precision and recall, rather than the arithmetic mean, it penalizes performance significantly if either precision or recall is low.
2.2. Notions of Model Fairness
Consider a machine learning classifier $h$, a set of features $X$, sensitive features $A\subset X$, and a set of labels $Y$. We can then define the following notions of fairness.
Definition 2.0 (Demographic Parity (Agarwal et al., 2018)).
The classifier $h$ satisfies Demographic Parity under a distribution over $(X,A,Y)$ if its prediction $h(X)$ is statistically independent of the sensitive feature $A$ i.e. $\mathbb{P}[h(X)=\hat{y}|A=a]=\mathbb{P}[h(X)=\hat{y}]$ for all $a,y$. For binary classification with $\hat{y}=\{0,1\}$, this is equivalent to $\mathbb{E}[h(X)|A=a]=\mathbb{E}[h(X)]$ for all $a$.
Definition 2.0 (Equalized Odds (Hardt et al., 2016; Agarwal et al., 2018)).
The classifier $h$ satisfies Equalized Odds under a distribution over $(X,A,Y)$ if its prediction $h(X)$ is conditionally independent of the sensitive feature $A$ given the label $Y$ i.e. $\mathbb{P}[h(X)=\hat{y}|A=a,Y=y]=\mathbb{P}[h(X)=\hat{y}|Y=y]$ for all $a$, $y$, and $\hat{y}$. For binary classification with $\hat{y}=\{0,1\}$, this is equivalent to $\mathbb{E}[h(X)|A=a,Y=y]=\mathbb{E}[h(X)|Y=y]$ for all $a$, $y$.
Equal Opportunity is a relaxed variant of Equalized Odds with $Y=1$ (Hardt et al., 2016). In other words, Equalized Odds require the true positive and false positive rates to be equal across all sensitive groups. In contrast, Equal Opportunity only requires the true positive rate to be equal across all sensitive groups.
Demographic Parity is the weakest notion of fairness, and Equalized Odds is the strongest. In our work, we use Equalized Odds as the fairness criterion. The disparity or unfairness for Equalized Odds is the Equalized Odds Difference (EOD), defined as the maximum absolute difference between the true and false positive rates across the sensitive groups. Mathematically, this is represented as
$EOD=\max\left(\begin{aligned} \lvert\mathbb{P}[h(X)=1|A=0,Y=1]-\mathbb{P}[h(X)=1|A=1,Y=1]\rvert,\\ \lvert\mathbb{P}[h(X)=1|A=0,Y=0]-\mathbb{P}[h(X)=1|A=1,Y=0]\rvert\end{aligned}\right)$
with sensitive features $A=\{0,1\}$.
3. Problem Statement
We consider pre-trained deep neural network (DNN) classifiers with the set of input variables $X$ partitioned into a protected set of variables (such as race, sex, and age) and non-protected variables (such as profession, income, and education). We further assume the output is a binary classifier that gives either favorable or unfavorable outcomes.
3.1. Syntax and Semantics of DNN
A deep neural network (DNN) encodes a function $\mathcal{D}:X\to[0,1]^{2}$ where $X$ consists of the set of protected attributes $X_{1}\times X_{2}\cdots\times X_{m}$ and non-protected attributes $X_{m+1}\cdots\times Y_{m+r}$. The DNN model is parameterized by the input dimension $m{+}r$, the output dimension $2$, the depth of hidden layers $n+1$, and the weights of its hidden layers $W_{0},W_{1},\ldots,W_{n}$. We describe the hidden layers with $\mathcal{M}\leftarrow[L_{I},L_{0},\ldots,L_{n},L_{O}]$, where $L_{I}$ and $L_{O}$ are the input and output layers, respectively, and ${L_{i}},\forall i\in[0,n]$, is the hidden layers. We assume that there exists a subset of neurons $\mathcal{N}\in{L_{i}},\forall i\in[0,n]$ in the hidden layers that disparately contribute to unfairness.
Let $L_{i}$ be the output of layer $i$ that implements an affine mapping from the output of previous layer $L_{i-1}$ and its weights $W_{i-1}$ for $1\leq i\leq n$ followed by a fixed non-linear activation unit (e.g., ReLU defined as $L_{i-1}\mapsto\max\left\{W_{i-1}.L_{i-1},0\right\}$) for $1\leq i\leq n$. Let $L_{i}^{j}$ be the output of neuron $j$ at layer $i$ that is $L_{i}^{j}(x)=\texttt{ReLU}\left(\sum_{j=1}^{|L_{i}|}w_{i}L_{i-1}^{j}\right)$. The output is the likelihood of favorable and unfavorable outcomes. The predicted label is the index of the maximum likelihood, $\mathcal{D}(x)=\max_{i}L_{O}(x)(i)$.
3.2. Inference Time Dropout for Fairness
Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of neural networks by preventing overfitting. The key idea is that randomly dropping out neurons in a layer during training can help prevent overfitting of the neural network. Dropout randomly sets all $w_{i}$ to 0 for each neuron in the hidden layers with probability $p$ during training. Once training is complete, dropout is not used, and all the neurons in the network are utilized to make predictions. While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the model during inference after training can significantly improve fairness with a minimal impact on model performance. However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
We consider a binary vector as the neuron state with $s=\{0,1\}^{N}$, where $N=\sum_{i=0}^{n}|L_{i}|$, $s_{i}$ indicates whether the neuron $i$ is dropped or not and $n$ is the number of layers. A pre-trained DNN model $\mathcal{D}$ does not include any dropouts, and hence all the indicators are 00.
Definition 3.0 (Desirable Dropout of Fairness vs. Utility).
Given a DNN model $\mathcal{D}$ trained over a dataset $X$; the search problem is to infer a repaired DNN model $\mathcal{D}^{\prime}$ by dropping a subset of neurons $I$ in the binary neuron state, i.e., $s_{i}=1$ for any $i\in I$, such that (1) the model bias (e.g., EOD) is maximally reduced, (2) the model performance (e.g., F1-score) is minimally degraded, and (3) the model structure in terms of the numbers of inputs, outputs, and hidden layers remains the same as compared to the original pre-trained model $\mathcal{D}$.
A brute-force search would be exponential in the size of DNN as we have $2^{|N|}$ possible subsets. The running time becomes prohibitively expensive, even for small neural networks. We define desirable states as those states that have a good fairness-performance tradeoff. To find the desirable subset of neurons, we explore different types of randomized algorithms to improve fairness via model inference dropout.
4. Approach
Our approach comprises two randomized search algorithms, namely Simulated Annealing (SA) and Random Walk (RW).
Figure 1. Overview of the NeuFair: a post-processing bias mitigation approach.
4.1. Simulated Annealing Search
We formulate the problem of finding the desirable subset of neurons as a discrete optimization problem and solve it using Simulated Annealing. Simulated Annealing (SA) (Bertsimas and Tsitsiklis, 1993) is a probabilistic algorithm that finds the global minima (w.r.t some cost function) in a large search space with high probability. Algorithm 1 presents a generic template to apply SA to a search space optimization problem. Figure 1 overviews the steps in our bias mitigation approach with the simulated annealing search. We now define the core concepts used in our SA algorithm.
1
Input: Search space $\mathcal{S}$, neighborhood relation $\Gamma$, cost function $c$, initial temperature $T_{0}$, initial state $s_{0}$
Output: Best solution found.
2
[…]
4
5while ‘stopping criterion’ is not met do
6       $s^{\prime}\leftarrow\text{Generate}(s)$
7      Compute $\Delta E=\text{cost}(s^{\prime})-\text{cost}(s)$
8      if $\Delta E\leq 0$ then
9             Accept transition i.e. $s\leftarrow s^{\prime}$
10             Update best solution found i.e. $s^{*}\leftarrow s^{\prime}$
11      else
12             if $e^{-\Delta E/T_{m}}\geq\text{UniformSample}(0,1)$ then
13                   Accept transition i.e. $s\leftarrow s^{\prime}$
14
15
16      $T_{m+1}\leftarrow\text{Update}(T_{m})$
17       $m\leftarrow m+1$
18
return $s^{*}$
Algorithm 1 Generic Simulated Annealing (SA) Procedure
Search Space
For a DNN $\mathcal{M}\leftarrow[L_{I},L_{0},\ldots,L_{n},L_{O}]$ we define a state of our search space $\mathcal{S}$ as a binary sequence $s\in\{0,1\}^{N}$, where $N=\sum_{i=0}^{n}|L_{i}|$ [^1] For memory reasons, we store a state as a decimal number rather than a $N$-bit binary sequence, where the binary sequence $s$ is mapped to the unique decimal number which has binary expansion $1s$. . The $i$th element of $s$, denoted $s_{i}$ is $1$ if neuron $i$ is dropped and 00 otherwise [^2] The neurons are numbered according to some total order; the choice of total order is insignificant as long as we fix the mapping of a neuron to its position in the binary sequence.. For example, consider a DNN $\mathcal{M}\leftarrow[L_{I},L_{0},L_{1},L_{O}]$, where $|L_{0}|=3$ and $|L_{1}|=3$. Then, the state $s$ given by:
$s=\underbrace{(0,1,0}_{L_{0}},\underbrace{0,0,1)}_{L_{1}}$
drops the second neuron in the first layer and the third neuron in the second layer.
Instead of allowing our search space $\mathcal{S}$ to drop every possible subset of neurons possible (which would ensure $\mathcal{S}$ has size $2^{N}$), we restrict the size of $\mathcal{S}$ by fixing an upper and lower bound on the number of neurons that can be dropped from the DNN. Let $n_{l}$ and $n_{u}$ ($n_{l}\leq n_{u}$) denote the minimum and maximum number of neurons allowed by our DNN, respectively. We can then formally define $\mathcal{S}$ as:
(1) $\mathcal{S}:=\{s\in\{0,1\}^{N}\;|\;n_{l}\leq HW(s)\leq n_{u}\}$
where $HW(s)$ denotes the hamming weight of $s$. Restricting our search space with a conservative estimate of the lower and upper bound would have a minimal impact on our ability to find the desirable subset, as dropping too many neurons will reduce the model utility to less than acceptable levels, whereas dropping too few neurons will improve fairness by only a marginal amount. However, asymptotically, a conservative estimate of the lower and upper bound still generates prohibitively large search spaces. With the bounds $n_{l}$ and $n_{u}$, the cardinality of the search space $|\mathcal{S}|=\sum_{i=n_{l}}^{n_{u}}{N\choose i}=\Omega(\frac{2^{N}}{\sqrt{N}})$ if $n_{l}<\frac{N}{2}<n_{u}$, which rules out brute-force as a viable option.
Neighborhood Relation and ‘Generate’ subroutine.
The neighborhood of any state $s\in\mathcal{S}$, denoted $\Gamma(s)$, is defined as the set of all states that are at a hamming distance of 1. Mathematically, this is defined as
$\Gamma(s):=\{s^{\prime}\in\mathcal{S}\;|\;HD(s,s^{\prime})=1\}$
where $HD(s,s^{\prime})$ denotes the hamming distance between $s$ and $s^{\prime}$. With our definition of the search space and neighborhood of a state, the entire search space graph can be viewed as a subset of the $N$-dimensional hypercube, where the vertices of the hypercube represent the states of our search space, and the edges of the hypercube represent the neighborhood relation. The Generate subroutine on input $s\in\mathcal{S}$ uniformly samples a neighbor $s^{\prime}$ from $\Gamma(s)$ and returns $s^{\prime}$. This is equivalent to uniformly sampling an index position $i$ from $[1,n]$ and subsequently uniformly flipping bit $s_{i}$ in the binary sequence $s$ to get $s^{\prime}$. We make the following observation on our underlying search space graph [^3] The search space graph is a graph with $\mathcal{S}$ as the set of vertices and edges defined according to the neighborhood relation $\Gamma$ i.e. $(s,s^{\prime})$ is an edge iff $s^{\prime}\in\Gamma(s)$. As our neighborhood relation is symmetric, our search space graph is undirected.:
Lemma 4.1.
If $n_{l}<n_{u}$, then our search space graph is connected and has a diameter less than $N$. Moreover, the distance between any two states $s,s^{\prime}\in\mathcal{S}$ is given by $HD(s,s^{\prime})$.
5. Experiments
We pose the following research questions:
  1. RQ1
    How successful are randomized algorithms in fairness repairing of DNNs via dropouts?
  2. RQ2
    Are there dropout strategies that improve fairness and utility together?
  3. RQ3
    What are the design considerations of search algorithms for efficient and effective fairness repair of DNNs via dropout?
  4. RQ4
    How do dropout strategies compare to the state-of-the-art post-processing (bias) mitigators?
5.1. Datasets and Models
We evaluate NeuFair with five different datasets from fairness literature. For two of the datasets, we consider two different protected groups, which effectively results in a total of 7 different benchmarks. Table 5.3 presents of an overview of the datasets and protected groups. The Adult Census Income (Dua and Graff, 2017b), Bank Marketing (Dua and Graff, 2017a), Compas Software (ProPublica, 2021), Default Credit  (Dua and Graff, 2017c), and Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) are binary classification tasks to predict whether an individual has income over $50$K, is likely to subscribe, has a low reoffending risk, is likely to default on the credit card payment, and is likely to utilize medical benefits, respectively.
We used DNNs as the machine learning model, trained over the dataset benchmarks. Table 1 highlights the architectures of the DNN models used for each dataset. We use ReLU as the activation function after the linear layers. For the Bank, Default, and Compas datasets we use a dropout of 0.2 during training. For the Adult and MEPS16 datasets, we set the dropout to 0.1. For the Compas dataset, we use Adam as the optimizer with a learning rate of $0.001$. For the other datasets, we use SGD with a learning rate of $0.01$. For data preprocessing, we utilize standard techniques such as one-hot encoding for categorical features followed by min-max or standard scaling for numerical features. For the Compas dataset, we use a version of the dataset that has 12 features after feature selection that was used in (Tizpaz-Niari et al., 2022). The modified Compas dataset is available online [^4] https://github.com/Tizpaz/Parfait-ML/blob/main/subjects/datasets/compas.
Table 1. Architectures of the DNNs used for the datasets.
Dataset
Model Architecture
$[L_{I},L_{0},\ldots,L_{n},L_{O}]$
Adult Census Income [34, 64, 128, 64, 1]
Compas Software [12, 32, 32, 1]
Bank Marketing [32, 32, 32, 1]
Default Credit [30, 16, 16, 16, 1]
Medical Expenditure (MEPS16) [138, 128, 128, 128, 1]
5.2. Technical Details
We run all our experiments on a desktop running Ubuntu 22.04.3 LTS with an Intel(R) Core(TM) i7-7700 CPU @ 3.60GHz processor, 32GB RAM, and a 1 TB HDD. The neural networks and repair algorithms were implemented using python3.8, torch==2.0.1,
numpy==1.24.3, and scikit-learn==1.3.1.
5.3. Experimental Setup
For each dataset, we evaluate the performance of the NeuFair algorithms using $10$ different random seeds. The seeds are set for torch, numpy, and scikit-learn libraries before DNN training. Specifically, the seed determines the randomness of the training, validation, and test splits and the randomness of model training, such as sampling batches and initializing the weights of the DNN. The NeuFair algorithms themselves are not seeded. Each SA and RW run is unique.
During training, the training dataset was used for gradient descent, and the validation dataset was used to monitor the F1 score and determine the hyperparameters. We use the model from the epoch with the highest validation dataset. The same validation dataset is used by NeuFair to identify the desirable set of neurons to drop. Finally, we evaluate the fairness of the original and fixed (repaired) models using the test dataset.
Unless otherwise specified, all runs reported in the paper have a time-out limit of $1$ hour with the following cost function:
$C(s)=EOD_{s}+3.0\cdot EOD_{s_{0}}\cdot\mathbbm{1}(F1_{s}<(0.98\cdot F1_{s_{0}})).$
The F1 threshold is 98% of the validation F1 score of the unfair neural network, i.e., we tolerate a 2% degradation in the F1 score to improve fairness. A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold. For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons depending on the overall size of the network. The $n_{u}$ values are 50, 24, 24, 20, and 135 for Adult, Compas, Bank, Default, and MEPS16 respectively. For all datasets, we use a train/validation/test split of 60%/20%/20%.
Table 2. Datasets used in our experiments
{tabu}
—l—l—l—ll—ll— Dataset —Instances— —Features— Protected Groups   Outcome Label
Group1 Group2 Label 1 Label 0
Adult Census (Dua and Graff, 2017b) $48,842$ $14$ Sex-Male Sex-Female High Income Low Income
Income Race-White Race-Non White
Compas Software (ProPublica, 2021) $7,214$ $28$ Sex-Male Sex-Female Did not Reoffend Reoffend
Race-Caucasian Race-Non Caucasian
Bank Marketing (Dua and Graff, 2017a) $45,211$ $17$ Age-Young Age-Old Subscriber Non-subscriber
Default Credit (Dua and Graff, 2017c) $13,636$ $23$ Sex-Male Sex-Female Default Not Default
Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) $15,675$ $138$ Race-White Race-Non White Utilized Benefits Not Utilized Benefits
Table 3. Fairness and model utility metrics of the unrepaired (original) neural network.
Dataset Unfair Neural Network
Validation Test
EOD F1 Accuracy EOD F1 Accuracy
Adult (Sex) $9.704\%\pm 1.405$ $0.68\pm 0.005$ $0.857\pm 0.003$ $11.639\%\pm 2.326$ $0.667\pm 0.008$ $0.851\pm 0.003$
Adult (Race) $8.352\%\pm 2.806$ $8.251\%\pm 3.195$
COMPAS (Sex) $2.307\%\pm 0.679$ $0.968\pm 0.002$ $0.97\pm 0.002$ $2.522\%\pm 0.817$ $0.967\pm 0.004$ $0.969\pm 0.004$
COMPAS (Race) $2.022\%\pm 0.884$ $2.96\%\pm 1.088$
Bank $13.752\%\pm 2.567$ $0.553\pm 0.006$ $0.842\pm 0.003$ $14.665\%\pm 2.114$ $0.553\pm 0.004$ $0.84\pm 0.003$
Default $9.068\%\pm 1.782$ $0.538\pm 0.006$ $0.774\pm 0.007$ $8.962\%\pm 1.772$ $0.53\pm 0.006$ $0.769\pm 0.007$
MEPS16 $20.167\%\pm 2.32$ $0.543\pm 0.009$ $0.79\pm 0.005$ $20.641\%\pm 2.527$ $0.533\pm 0.01$ $0.788\pm 0.009$
Table 4. Fairness and model utility metrics of the repaired neural network.
Dataset Repaired Neural Network
Simulated Annealing Random Walk
Validation Test Validation Test
EOD F1 Accuracy EOD F1 Accuracy EOD F1 Accuracy EOD F1 Accuracy
Adult (Sex) $5.618\%\pm 0.807$ $0.668\pm 0.005$ $0.856\pm 0.004$ $7.259\%\pm 1.697$ $0.652\pm 0.01$ $0.849\pm 0.004$ $6.01\%\pm 0.704$ $0.667\pm 0.005$ $0.855\pm 0.003$ $7.358\%\pm 1.063$ $0.652\pm 0.01$ $0.849\pm 0.004$
Adult (Race) $3.298\%\pm 2.069$ $0.667\pm 0.005$ $0.854\pm 0.003$ $4.976\%\pm 1.816$ $0.656\pm 0.008$ $0.849\pm 0.004$ $3.965\%\pm 1.665$ $0.667\pm 0.005$ $0.853\pm 0.003$ $4.785\%\pm 2.085$ $0.658\pm 0.009$ $0.849\pm 0.003$
COMPAS (Sex) $0.468\%\pm 0.548$ $0.955\pm 0.004$ $0.959\pm 0.003$ $2.921\%\pm 1.446$ $0.954\pm 0.08$ $0.957\pm 0.008$ $0.54\%\pm 0.535$ $0.956\pm 0.005$ $0.959\pm 0.004$ $2.233\%\pm 1.022$ $0.954\pm 0.007$ $0.957\pm 0.006$
COMPAS (Race) $0.56\%\pm 0.71$ $0.955\pm 0.002$ $0.959\pm 0.002$ $2.239\%\pm 1.003$ $0.954\pm 0.005$ $0.957\pm 0.004$ $0.58\%\pm 0.715$ $0.956\pm 0.003$ $0.959\pm 0.002$ $2.159\%\pm 1.07$ $0.955\pm 0.004$ $0.958\pm 0.004$
Bank $0.871\%\pm 0.509$ $0.547\pm 0.007$ $0.892\pm 0.01$ $7.257\%\pm 3.533$ $0.537\pm 0.014$ $0.888\pm 0.01$ $1.714\%\pm 0.921$ $0.548\pm 0.006$ $0.882\pm 0.01$ $7.595\%\pm 2.733$ $0.548\pm 0.008$ $0.881\pm 0.008$
Default $1.14\%\pm 0.98$ $0.529\pm 0.007$ $0.794\pm 0.016$ $2.749\%\pm 0.827$ $0.519\pm 0.006$ $0.79\pm 0.015$ $2.045\%\pm 1.251$ $0.53\pm 0.007$ $0.794\pm 0.015$ $3.124\%\pm 0.937$ $0.523\pm 0.005$ $0.79\pm 0.015$
MEPS16 $4.589\%\pm 1.294$ $0.535\pm 0.01$ $0.86\pm 0.006$ $8.426\%\pm 2.311$ $0.507\pm 0.02$ $0.853\pm 0.005$ $6.622\%\pm 1.183$ $0.535\pm 0.01$ $0.856\pm 0.005$ $9.86\%\pm 2.623$ $0.513\pm 0.018$ $0.851\pm 0.007$
Figure 2. Evolution of the cost of the best state (logarithmic scale) found during the search on validation dataset using Simulated Annealing (SA) with 7 benchmarks and 10 seeds each.
Figure 3. Evolution of the cost of the best state (logarithmic scale) found during the search on validation dataset using Random Walk (RW) with 7 benchmarks and 10 seeds each.
5.3.1. Effectiveness of randomized algorithms on mitigating unfairness using dropout (RQ1)
The first research question is answered by analyzing the data in Figure 2, Figure 3, Table 3, and Table 4.
Figure 2 and Figure 3 highlight the cost of the best state found for the SA and RW over time, respectively, for 10 different seeds per data subject using the logarithmic scale. In all cases, we see a positive trend where the cost improves over time, and the search identifies better subsets of neurons to drop. We note the seemingly flat plots correspond to those cases where a good state was found in the beginning. The exception is the black line in the sub-plot for the Default dataset in both figures as it corresponds to a failed run; i.e., the best state found has an F1 score lesser than 0.98×0.98$\times$0.98 × the F1 score of the initial state. Out of 70 runs (10 per benchmark), the SA and RW found a desirable state with 0.98×0.98$\times$0.98 × the initial F1 score in 69 cases. In addition, we re-run the experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$. The progress graphs of the RW strategy are “steep” initially and rapidly improve as the initial states are unfair and also heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold. However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability. So, the progress graphs are not as steep initially.
Table 3 and Table 4 highlight the mean and $95\%$ confidence intervals for the F1 score, accuracy, and EOD for all data subjects aggregated across 10 seeds. We observe that in all cases, both NeuFair algorithms improve the EOD. However, the SA algorithm outperformed RW in terms of EOD in all cases for the validation set while ensuring that the validation F1 score reduces by at most $2\%$. We can conclude that SA is the more effective algorithm in identifying the desirable state during the search procedure, subject to the F1 threshold. On the test set, we observe that both NeuFair algorithms improve the EOD. SA outperforms RW for 4 datasets (Adult (Sex), Bank, Default, and MEPS16). We observe the largest absolute improvement of $12.215\%$ in EOD ($20.641\%$ to $8.426\%$ ) for the MEPS16 dataset using the SA algorithm. The largest relative improvement of $69.43\%$ ($8.962\%$ to $2.749\%$) was observed in the Default dataset using SA. However, SA did not improve the EOD on the COMPAS (Sex), which is a pathological case.
Table 5. Fairness metrics for COMPAS (Sex)
Model EOD for 10 seeds
Unfair 2.2% 2.5% 2.35% 1.87% 4.1% 2.15% 1.38% 4.26% 0.79% 3.6%
Repaired (SA) 1.97% 4.28% 2.35% 4.58% 4.37% 0.12% 0.47% 4.83% 5.5% 0.75%
Repaired (RW) 1.25% 3.61% 3.19% 2.5% 2.4% 0.53% 0.28% 4.54% 3.07% 0.92%
Pathological Case. Table 5 highlights the EOD scores for COMPAS (Sex) with the original and repaired networks. Here, we observe that SA improves the EOD in 4 cases, does not improve in 1 case, and has a worse EOD in 5 cases. RW improves the EOD in 5 cases and has a worse EOD in the other 5 cases. There are two reasons for the different trends in COMPAS (Sex). First, there is a discrepancy in the original model statistics between the validation and test splits. The difference arises because the DNN tends to overfit slightly on the training/validation sets. This is evident from model metrics in Table 3. Second, the COMPAS datasets by default are quite fair (according to EOD) compared to other datasets as their unfairness is around 2%. During the search, the SA algorithm correctly identifies a desirable state according to the validation dataset and performs better than RW. However, given the low initial unfairness and the model’s tendency to perform better on the validation/training sets, we observe minor discrepancies in the test set. The average unfairness, however, increases by 0.399% (2.522% to 2.921%), which is negligible. For both NeuFair algorithms, the validation fairness improves in the COMPAS (Sex) experiment.
Answer RQ1: The randomized algorithms effectively mitigate unfairness via dropouts by de-activating a desirable subset of neurons. On the test set, Fairness improves by up to 69%. The SA algorithm performs better than the RW algorithm.
5.3.2. Dropout strategies improving both fairness and utility (RQ2)
The second research question is answered by observing the F1 score and accuracy (model utility metrics), besides the EOD score (fairness) in Table 3 vs. Table 4. For all datasets, the F1 score decreases for the validation and test sets. The decrease in F1 score is accompanied by an improvement in fairness which highlights the tradeoff between the model utility and fairness. However, we observe that the accuracy increases for the Bank, Default, and MEPS16 datasets. The biggest improvement in accuracy is observed in MEPS16, where the Validation and Test accuracy increase from 0.79 and 0.788 to 0.86 and 0.853 for SA and to 0.856 to 0.851 for RW. For Default, the validation and test accuracies increase from 0.774 and 0.769 to 0.794 and 0.79 for both SA and RW. For Bank, the validation and test accuracies increase from 0.842 and 0.84 to 0.892 and 0.88 for SA and 0.882 and 0.881 for RW.
The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is negative class and 1 is positive) as more and more neurons are dropped out from the DNN. The increase in negative predictions favors the true negatives as the datasets are higly imbalanced with more negative samples. The Default, Bank, and MEPS16 datasets have 78%, 88%, and 83% of the data belonging to the negative class. The trained DNN models were optimized for F1 score on the validation dataset to improve performance on the underrepresented positive class. The accuracy increases as the increase in true negatives outweighs the drop in true positives. The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and false negatives increase.
Answer RQ2: The overall accuracy, as a model utility metric, may improve along with fairness using dropout strategies, as accuracy does not account for false positives and false negatives in imbalanced datasets. However, the F1 score model utility metric decreases as fairness increases.
5.3.3. Hyperparameters of randomized algorithms and fairness (RQ3)
The NeuFair algorithms have 4 hyperparameters: F1 threshold multiplier, the minimum and maximum number of neurons to drop, time-out limit, and F1 penalty multiplier.
  • The F1 threshold multiplier is inversely proportional to fairness improvement. For example, as highlighted in RQ1, the fairness of the Default experiment improves when the F1 threshold is reduced from 0.98 to 0.96.
  • Decreasing the minimum number of neurons and increasing the maximum number of neurons can have a positive effect on fairness, provided the time-out limit increases. Consider two ranges $[n_{l1},n_{u1}]$ and $[n_{l2},n_{u2}]$ such that $n_{l1}<n_{l2}<n_{u2}<n_{u1}$. The search space of $[n_{l2},n_{u2}]$ is then a sub-space of $[n_{l1},n_{u1}]$.
  • The time-out positively affects fairness as the NeuFair algorithms have more time to explore the search space.
  • The F1 penalty multiplier has a more nuanced effect on the unfairness as it controls the exploration vs. exploitation trade-offs. A low penalty multiplier increases the probability of state transitions when the F1 score is less than our threshold, increasing the SA run’s randomness to explore more of the search space outside of the current best state. A high penalty multiplier would decrease the probability of state transitions, thereby encouraging the SA algorithm to exploit the search space near the current best state. The penalty multiplier ‘p’ directly affects the cost difference $\Delta E$ between states with acceptable and unacceptable F1 scores. From Line 14 in Algorithm LABEL:alg:neuron_repair, we can see that a higher cost difference results in a lower transition probability. For a fixed temperature, a higher ‘p’ would increase the cost difference between good and bad states, thereby decreasing the probability of transitioning and encouraging exploitation. A lower ‘p’ would result in a lower cost difference and increase the transition probability, thereby encouraging exploration. To determine the effect of the penalty multiplier, we run experiments with the Adult, MEPS16, Default, and Bank datasets with one seed by varying the F1 penalty multiplier $p\in\{0.5,1.0,1.5,2.0,2.5,3.0\}$ and F1 threshold multiplier to 0.98. Figure 4 shows the results. The red dot in the Default dataset corresponds to a run in which SA failed to find a desirable state within the F1 threshold. We can see that Adult dataset performs better on lower F1 penalties. The Bank dataset has a “sweet” spot in between while Default performs worse on intermediate F1 penalties. The MEPS16 does not follow a clear trend. The results show that the F1 penalty multiplier is a hyperparameter that trades off explorations vs. exploitation and requires tuning for each dataset and model.
Figure 4. Effect of F1 penalty multiplier on the EOD
Answer RQ3: The fairness of the repaired model can improve as the F1 threshold decreases, the number of neurons to drop increases, and the time-out limit increases for both NeuFair-SA and NeuFair-RW algorithms. The F1 penalty multiplier controls a trade-off between explorations vs. exploitation, which requires tuning for each specific benchmark. A higher F1 penalty multiplier encourages SA to exploit the local solutions, whereas a lower penalty encourages exploring the global space.
5.3.4. Comparing to the state-of-the-art (RQ4)
We compare the efficacy of NeuFair against a state-of-the-art post-processing bias mitigator. Dice (Monjezi et al., 2023) used a fault localization technique via do logic to pinpoint a single neuron that significantly influences fairness. The approach computes the amounts of variation in the outcomes of each layer that depends on the sensitive attributes and de-activates a single neuron with the highest effects on fairness. Since the search space evaluates a single neuron, we brute-force the entire search space of one neuron dropout in a linear time. We note that the best improvement, feasible by Dice, cannot be more than the brute-force search. Table 6 highlights the test fairness after dropping a single neuron that most impacts fairness. The results show that NeuFair substantially outperforms DICE in improving fairness. We also note a similar trend with COMPAS (Sex), where the unfairness increases slightly. The largest discrepancy in fairness improvement between NeuFair and Dice is observed in the MEPS16 dataset, with an absolute difference of 10.778% and a relative improvement of 56.124% (19.204% and 8.426%) using the SA algorithm.
Answer RQ4: NeuFair outperforms the state-of-the-art unfairness mitigation algorithm Dice (Monjezi et al., 2023) across all benchmarks. The relative improvement of NeuFair is 56% higher than Dice.
Table 6. Fairness and Utility of Dice (Monjezi et al., 2023).
Dataset Test Dataset
EOD F1 Accuracy
Adult (Sex) $10.453\%\pm 2.266$ $0.658\pm 0.007$ $0.851\pm 0.003$
Adult (Race) $8.092\%\pm 3.253$ $0.662\pm 0.008$ $0.851\pm 0.003$
COMPAS (Sex) $3.229\%\pm 0.774$ $0.964\pm 0.005$ $0.957\pm 0.004$
COMPAS (Race) $2.964\%\pm 1.088$ $0.965\pm 0.003$ $0.968\pm 0.003$
Bank $12.205\%\pm 2.731$ $0.567\pm 0.005$ $0.862\pm 0.01$
Default $5.845\%\pm 1.816$ $0.528\pm 0.004$ $0.791\pm 0.007$
MEPS16 $19.204\%\pm 2.592$ $0.54\pm 0.012$ $0.8\pm 0.009$
6. Discussion
Table 7. Simulated Annealing compared to Brute-force search strategies.
Seed Unfair DNN EOD Repaired Neural Network EOD
SA Brute Force EOD Delta
1 14.087% 6.441% 6.415% 0.026
2 8.874% 5.306% 5.306% 0.0
3 8.352% 5.755% 5.755% 0.0
4 17.896% 13.823% 13.743% 0.08
5 8.062% 5.304% 5.25% 0.054
Table 8. Distribution of Best, Good, and Bad states in the search space.
Seed Best State Good State Bad State
Count Likelihood Count Likelihood Count Likelihood
1 1 $2.3\times 10^{-8}$ 43,789 0.001 41,669,509 0.967
2 11 $2.5\times 10^{-7}$ 6,717 $1.0\times 10^{-4}$ 40,062,436 0.93
3 8 $1.8\times 10^{-7}$ 1,795 $4.1\times 10^{-5}$ 42,997,093 0.998
4 1 $2.3\times 10^{-8}$ 790 $1.8\times 10^{-5}$ 42,467,272 0.985
5 1 $2.3\times 10^{-8}$ 13,623 $3.16\times 10^{-4}$ 40,197,630 0.933
Effectiveness of Randomized Algorithms. To understand why randomized algorithms are effective in mitigating unfairness, we compare them to brute-force strategies. We create a small DNN for the Adult (Sex) dataset with 2 hidden layers, each with 16 neurons, to allow the brute force search to explore the entire space. We use a restricted search space with $n_{l}=4$ and $n_{u}=9$ that has a total of $43,076,484$ states. The time-out for SA runs is one hour. The brute force runs take 60 hours to explore the entire search space. Table 7 highlights the results of both search strategies. SA finds the global optimal solution for two seeds in the search space. The EOD Delta column highlights the difference in cost of the states identified by SA and brute force. The highest cost difference of $0.054$ shows the efficacy of randomized algorithms in identifying good states. SA is faster in reaching a good state compared to brute force as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost. Another reason for SA’s effectiveness is that it uses a probabilistic transition function that can transition to a state that has a worse cost and escape local minima. Gradient-based methods often struggle with the problem of getting stuck in local minima. Table 8 shows the distribution of the Best, Good, and Bad states. The best states are the states with the global optimal cost. The cost difference between the good and best states is less than 0.05. Bad states are states whose F1 scores are less than the 98% of the baseline F1 score. We observe that the best and good states are a negligible amount of the search space. On the other hand, over 90% of the states are bad states. This suggests that the good states occupy small “pockets” of the search space. The cost gradient leads SA to effectively identify such pockets. The probabilistic transition function enables SA to explore the region without getting stuck in local minima.
Limitations. Since we used randomized algorithms, our results might be sub-optimal. While simulated annealing provides statistical guarantees on the confidence and running time to reach the optimal state, in practice such guarantees might require multiple runs of algorithms over a long period of time. Additionally, we observe a discrepancy between the validation and test fairness improvements where the test set improves less than the validation set in both NeuFair algorithms. For our experiments, we use Equalized Odds Difference (EOD) only. However, the cost function can be modified accordingly with any fairness metric.
Threat to Validity. To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines and repeat the experiments 10 times, reporting both the average and 95% confidence intervals. In addition, we study the progress of algorithms during their runs, not just the final outcomes. To ensure that our results are generalizable and address external validity, we perform our experiments on seven DNN benchmarks with various architectures. While similar architectures have been used in the prior works (Monjezi et al., 2023; Zhang et al., 2020b; Zheng et al., 2022; Zhang et al., 2021), it is an open problem whether these datasets and DNN models are sufficiently representative for showing the effectiveness of NeuFair.
7. Related Work
Since the main focus of this paper is unfairness mitigation, we primarily focus on the prior work on mitigation.
A) Pre-processing Techniques exploit the space of input data to mitigate fairness defects (Galhotra et al., 2022; Feldman et al., 2015; Kamiran and Calders, 2012; Zemel et al., 2013). Reweighting (Kamiran and Calders, 2012) augments the dataset where the datapoints in each group-label combination are weighted differently to ensure fairness. The reweighting process minimizes the discrepancy between the observed and expected probabilities of the favorable outcome occurring with the discriminated sensitive attribute. FairMask (Peng et al., 2022) uses the fact that protected attributes can be inferred by some combinations of other non-protected attributes (the “proxy” problem) to reduce the influence of protected attributes in the inference stage. Fair-SMOTE (Chakraborty et al., 2021a) used under-sampling and over-sampling techniques (Chawla et al., 2002) to balance data based on class and sensitive attributes. MAAT (Chen et al., 2022) proposes an ensemble training approach to mitigating unfairness. They propose a framework to train fairness and performance ML models and combine their outputs to make predictions. The performance model is identical to traditional ML algorithms and remains unchanged. For the fairness model, they design a debugging strategy based on prior work (Wick et al., 2019; Chakraborty et al., 2021b) that attribute bias in training data to selection bias and label bias. LTDD (Li et al., 2022) hypothesizes that biased features in training data contributes to unfairness. To obtain unbiased features, they develop a linear regression based algorithm that measures the association between non-sensitive and sensitive features in the dataset. The non-sensitive features are then modified to remove the biased parts to become independent of the sensitive features. After debiasing, the sensitive features are dropped from the dataset.
The aforementioned works tamper with the precious training data samples that might lead to unrealistic representations of different protected groups. Additionally, it may not always be feasible to modify training data. Our work is tangential as it focuses on post-processing mitigation that does not modify or access the training dataset and works on pre-trained unfair DNNs.
B) In-processing techniques aim to improve fairness during training. One method, known as adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions. Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the learning objective (loss function), aiming to balance both accuracy and fairness. These strategies necessitate alterations in either the loss function or the parameters of the ML models. Seldonian (Thomas et al., 2019; Metevier et al., 2019; Hoag et al., 2023) presented a technique that allows users to directly specify (arbitrary) undesirable behaviors as constraints and enforce them during training. Fairway (Chakraborty et al., 2020) combines pre-processing and in-processing mitigation techniques to improve fairness. Parfait-ML (Tizpaz-Niari et al., 2022) is a gray-box evolutionary search algorithm that explores the ML hyperparameters to find configurations that minimize fairness while maintaining an acceptable accuracy. While mitigating unfairness during training might be less intrusive than changing the training datasets; they often require changing training algorithms that might be only effective for a particular algorithm and fairness definition. In addition, the training process is often expensive and might not be feasible for existing systems.
C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022). Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post-hoc correction mechanism that modifies the output of a biased binary classifier. EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor. NeuFair, however, takes a fundamentally different approach to post-processing mitigation. First, NeuFair does not require training auxiliary models to mitigate unfairness. Second, NeuFair does not rely on randomization for fairness and generates a deterministic DNN classifier. Predictions during inference time are deterministic after the neurons are dropped from the DNN. Lastly, Woodworth et al. (2017) shows that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data. Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness. Care (Sun et al., 2022) presented Particle Swarm Optimisation (PSO) to repair DNNs. Faire (Li et al., 2023) proposes a method to repair unfair DNNs by altering the activations of neurons. Inspired by program repair techniques, Faire first groups neurons representing protected and non-protected features in the dataset into two categories. To create the two categories, Faire creates a clone of the unfair model and retrains it to predict the protected attribute. By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected. Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted. Compared to Faire, NeuFair does not retrain the model and does not separately use the protected attribute to identify unfair neurons. We find an unfair subset of neurons given the entire feature vector, thereby accounting for the causal influence of the protected features on the non-protected features in the training dataset. Our analysis considers a group of neurons together rather than identifying the contributions of individual neurons. FairNeuron (Gao et al., 2022) proposes an algorithm to selectively retrain neurons in a neural network to mitigate unfairness. On a high level, the algorithm identifies input-output neuron paths corresponding to each sample in the training dataset that induces bias in the prediction. Then, they split the dataset into biased and unbiased subsets corresponding to the neuron activations in the paths. Finally, the algorithm re-trains the model with both training data splits by enabling random dropouts on the neurons in the biased paths and disabling any dropouts for neurons in the un-biased paths. While FairNeuron utilizes dropouts to improve fairness, it fundamentally differs from NeuFair. First, we do not require access to the large training dataset, and we do not retrain the models. Second, NeuFair algorithms utilize deterministic dropout during inference time to mitigate unfairness with a flexible budget on the loss in F1 score, i.e., we identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly. Lastly, since we do not perform re-training, NeuFair does not update the weights of any neuron in the network.
8. Conclusion
In this paper, we tackle the problem of mitigating unfairness in pre-trained deep neural networks using the dropout method. We showed that the neural dropout problem over the DNN models is computationally hard and presented NeuFair, a family of randomized algorithms to efficiently and effectively improve the fairness of DNNs. Our experiments showed that NeuFair can identify an ideal subset of neurons to drop that disparately contribute to unfairness (leading to up to 69% fairness improvement) and outperform a state-of-the-art post-processing bias mitigator.
For future work, there are a few exciting directions. First, instead of choosing the best state found during the search, we can analyze the top $k$ states found during the run. Such a strategy could reduce the difference in improvement observed between the validation and test sets. We can extend to multi-valued protected attributes (age groups) or optimize for more than one protected attribute at a time. Second, we plan to leverage dropouts in the fine-tuning process to understand the effects of randomized algorithms on mitigating the unfairness and toxicity of large language models. Lastly, alternate strategies like genetic algorithms can be considered to solve the combinatorial optimization problem of neuron dropout. We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems. The primary difference between SA/RW and genetic algorithms is that genetic algorithms create a population of candidate solutions at each step instead of a single solution. Genetic algorithms can be promising as they enable us to explore more states of the search space simultaneously.
Data Availability. Our open-source tool NeuFair with all experimental subjects are publicly accessible: Link.

### v2

@@
 Abstract.
-This paper investigates the neural dropout method as a post-processing bias mitigation for deep neural networks (DNNs).
+This paper investigates neuron dropout as a post-processing bias mitigation for deep neural networks (DNNs).
 Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications.
-While neural networks are exceptionally good at finding statistical patterns from data, they are notorious for over-fitting the training datasets that may encode and amplify existing biases from the historical data.
-Existing bias mitigation algorithms often require either modifying the input dataset or modifying the learning algorithms.
-We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve fairness of pre-trained DNNs.
+While neural networks are exceptionally good at finding statistical patterns from data, they may encode and amplify existing biases from the historical data.
+Existing bias mitigation algorithms often require modifying the input dataset or the learning algorithms.
+We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve the fairness of pre-trained DNNs.
 However, finding the ideal set of neurons to drop is a combinatorial problem.
@@
 Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility.
-We show that our design of randomized algorithms are effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation.
+We show that our design of randomized algorithms is effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation.
 We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of search algorithms on the results.
@@
 Artificial intelligence (AI), increasingly deployed with deep neural network (DNN) components, has become an integral part of modern software solutions that assist in socio-economic and legal-critical decision-making processes such as releasing patients (Kharpal, 2018), identifying loan defaults (Group, 2018), and detecting tax evasion (Tankersley, 2023).
-Despite many advances made possible by AI, there are challenges that require understanding the dimensions and implications of deploying AI-driven software solutions.
+Despite many advances made possible by AI, some challenges require understanding the dimensions and implications of deploying AI-driven software solutions.
 One such concern about the trustworthiness of AI is discrimination.
@@
 Parole decision-making software was found to harm black and Hispanic defendants by falsely predicting a higher risk of recidivism than for non-Hispanic white defendants (Julia Angwin and Kirchne, 2021); Amazon’s hiring algorithm disproportionately rejected more female applicants than male applicants (Ingold and Soper, 2016); and data-driven auditing algorithm selected black taxpayers with earned income tax credit claims (EITC) at much higher rates than other racial groups for an audit (Tankersley, 2023).
-As evidenced by these examples, resulting software may particularly disadvantage minorities and protected groups and be found non-compliant with law such as the US Civil Rights Act (Blumrosen, 1978).
+As evidenced by these examples, resulting software may particularly disadvantage minorities and protected groups and be found non-compliant with laws such as the US Civil Rights Act (Blumrosen, 1978).
 Hence, helping programmers and users to mitigate unfairness in social-critical data-driven software systems is crucial to ensure inclusion in our modern, increasingly digital society.
@@
 Broadly, fairness mitigation can be applied in pre-processing stage (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing step (e.g., changing the loss function to include fairness constraints during training), or post-processing stage (e.g., changing the logic of a pre-trained model).
-However, when the decision logic of AI is encoded via deep neural networks, it becomes increasingly challenging to mitigate unfairness due to its black-box uninterpretable nature.
-Since the DNN-based software development is derived by principle of information bottleneck (Tishby et al., 2000), which provides a trade-off between over-fitting and parsimonious signals; we posit that over-fitted DNN models may contribute and amplify historical biases from the training datasets.
-Therefore, prevalent techniques such as dropout methods (Srivastava et al., 2014; Srivastava, 2013)—a process of randomly dropping neurons during training—might be effective in mitigating unfairness in the pre-trained DNN models.
-While the dropout strategy has been significantly used to prevent over-fitting during the training of DNNs, to the best of our knowledge, this is the first work to systematically leverage dropout methods for mitigating unfairness in pre-trained DNN models, as a post-processing mitigation method.
+However, when the decision logic of AI is encoded via deep neural networks, it becomes challenging to mitigate unfairness due to its black-box uninterpretable nature.
+We posit that a subset of neurons in a neural network disparately contributes to unfairness.
+Removing these neurons during inference as a post-processing operation on a trained DNN can improve fairness.
+Therefore, prevalent techniques such as dropout methods (Srivastava et al., 2014; Srivastava, 2013)—a process of randomly dropping neurons during training—might be effective in mitigating unfairness in pre-trained DNN models.
+While the dropout strategy has been significantly used to prevent over-fitting during the training of DNNs, to the best of our knowledge, this is the first work to systematically leverage dropout methods as a post-processing bias mitigation method for improving the fairness of pre-trained DNN models.
 However, finding the optimal subset of dropout neurons is a combinatorial problem that requires an exhaustive search in the subset of all combinations of neuron dropouts, and hence is intractable.
@@
 We design and implement simulated annealing (SA) and random walk (RW) strategies that efficiently explore the state-space of neuron dropouts where we encode the frontiers of fairness and utility in a cost function.
-We evaluate NeuFair over $7$ deep neural network benchmarks as trained over $5$ socially critical applications with significant fairness implications.
-We found that randomized strategies of neuron drops can successfully improve fairness (up to 69%) with a minimal utility degradation in the vast majority of cases.
+We evaluate NeuFair over $7$ deep neural network benchmarks trained over $5$ socially critical applications with significant fairness implications.
+We found that NeuFair can improve fairness (up to 69%) with minimal utility degradation in most cases.
 We also report a pathological case and reasons behind a failure of NeuFair.
-We also observe that NeuFair can improve fairness and utility together and provide intuitive explanations of such phenomena.
+We also observe that NeuFair can simultaneously improve fairness and utility and provide intuitive explanations of such phenomena.
 Furthermore, we examine different hyperparameter configuration options of randomized algorithms.
@@
 3.2. Inference Time Dropout for Fairness
-Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of neural networks by preventing overfitting.
-The key idea is that randomly dropping out neurons in a layer during training can help prevent overfitting of the neural network.
-Dropout randomly sets all $w_{i}$ to 0 for each neuron in the hidden layers with probability $p$ during training.
-Once training is complete, dropout is not used, and all the neurons in the network are utilized to make predictions.
-While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the model during inference after training can significantly improve fairness with a minimal impact on model performance.
+Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of DNNs by preventing overfitting.
+Dropout sets all $w_{i}$ to 0 for a random set of neurons in the hidden layers with some probability during training.
+Once training is complete, dropout is not used, and all the neurons in the DNN are utilized to make predictions.
+While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the DNN during inference after training can significantly improve fairness with a minimal impact on model performance.
 However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
@@
 Given a DNN model $\mathcal{D}$ trained over a dataset $X$; the search problem is to infer a repaired DNN model $\mathcal{D}^{\prime}$ by dropping a subset of neurons $I$ in the binary neuron state, i.e., $s_{i}=1$ for any $i\in I$, such that (1) the model bias (e.g., EOD) is maximally reduced, (2) the model performance (e.g., F1-score) is minimally degraded, and (3) the model structure in terms of the numbers of inputs, outputs, and hidden layers remains the same as compared to the original pre-trained model $\mathcal{D}$.
-A brute-force search would be exponential in the size of DNN as we have $2^{|N|}$ possible subsets.
-The running time becomes prohibitively expensive, even for small neural networks.
 We define desirable states as those states that have a good fairness-performance tradeoff.
+A brute-force search to find desirable neurons is exponential in the size of DNN as we have $2^{|N|}$ possible subsets.
+The running time becomes prohibitively expensive, even for small DNNs.
 To find the desirable subset of neurons, we explore different types of randomized algorithms to improve fairness via model inference dropout.
@@
 Algorithm 1 presents a generic template to apply SA to a search space optimization problem.
-Figure 1 overviews the steps in our bias mitigation approach with the simulated annealing search.
+Figure 1 overviews the steps in our bias mitigation approach with the SA search.
 We now define the core concepts used in our SA algorithm.
@@
 The Adult Census Income (Dua and Graff, 2017b), Bank Marketing (Dua and Graff, 2017a), Compas Software (ProPublica, 2021), Default Credit  (Dua and Graff, 2017c), and Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) are binary classification tasks to predict whether an individual has income over $50$K, is likely to subscribe, has a low reoffending risk, is likely to default on the credit card payment, and is likely to utilize medical benefits, respectively.
-We used DNNs as the machine learning model, trained over the dataset benchmarks.
+We used DNNs trained over the dataset benchmarks as the machine learning model.
 Table 1 highlights the architectures of the DNN models used for each dataset.
@@
 For each dataset, we evaluate the performance of the NeuFair algorithms using $10$ different random seeds.
-The seeds are set for torch, numpy, and scikit-learn libraries before DNN training.
+The seeds are set for the torch, numpy, and scikit-learn libraries before DNN training.
 Specifically, the seed determines the randomness of the training, validation, and test splits and the randomness of model training, such as sampling batches and initializing the weights of the DNN.
@@
 Each SA and RW run is unique.
-During training, the training dataset was used for gradient descent, and the validation dataset was used to monitor the F1 score and determine the hyperparameters.
-We use the model from the epoch with the highest validation dataset.
-The same validation dataset is used by NeuFair to identify the desirable set of neurons to drop.
-Finally, we evaluate the fairness of the original and fixed (repaired) models using the test dataset.
+During training, the training dataset was used for gradient descent, and the validation dataset was used for hyperparameter tuning.
+We use the model from the epoch with the highest validation F1 score.
+The validation dataset is also used by NeuFair to identify the desirable set of neurons to drop.
+We evaluate the fairness of the original and fixed (repaired) models using the test dataset.
 Unless otherwise specified, all runs reported in the paper have a time-out limit of $1$ hour with the following cost function:
 $C(s)=EOD_{s}+3.0\cdot EOD_{s_{0}}\cdot\mathbbm{1}(F1_{s}<(0.98\cdot F1_{s_{0}})).$
-The F1 threshold is 98% of the validation F1 score of the unfair neural network, i.e., we tolerate a 2% degradation in the F1 score to improve fairness.
+The F1 threshold is 98% of the validation F1 score of the unfair DNN, i.e., we tolerate a 2% degradation in the F1 score to improve fairness.
 A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold.
@@
 For all datasets, we use a train/validation/test split of 60%/20%/20%.
-Table 2. Datasets used in our experiments
+We set the threshold for the acceptance probabilities in the temperature initialization algorithm (Ben-Ameur, 2004) to $0.75$.
+Table 2. Datasets used in our experiments.
 {tabu}
@@
 Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) $15,675$ $138$ Race-White Race-Non White Utilized Benefits Not Utilized Benefits
-Table 3. Fairness and model utility metrics of the unrepaired (original) neural network.
+Table 3. Fairness and model utility metrics of the unrepaired (original) DNN.
 Dataset Unfair Neural Network
@@
 MEPS16 $20.167\%\pm 2.32$ $0.543\pm 0.009$ $0.79\pm 0.005$ $20.641\%\pm 2.527$ $0.533\pm 0.01$ $0.788\pm 0.009$
-Table 4. Fairness and model utility metrics of the repaired neural network.
+Table 4. Fairness and model utility metrics of the repaired DNN.
 Dataset Repaired Neural Network
@@
 In addition, we re-run the experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$.
-The progress graphs of the RW strategy are “steep” initially and rapidly improve as the initial states are unfair and also heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold.
+The progress graphs of the RW strategy are “steep” initially and as the initial states are unfair and are heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold.
 However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability.
@@
 We observe that in all cases, both NeuFair algorithms improve the EOD.
-However, the SA algorithm outperformed RW in terms of EOD in all cases for the validation set while ensuring that the validation F1 score reduces by at most $2\%$.
+However, SA outperformed RW in all cases for the validation set while ensuring that the F1 score reduces by at most $2\%$.
 We can conclude that SA is the more effective algorithm in identifying the desirable state during the search procedure, subject to the F1 threshold.
@@
 SA outperforms RW for 4 datasets (Adult (Sex), Bank, Default, and MEPS16).
-We observe the largest absolute improvement of $12.215\%$ in EOD ($20.641\%$ to $8.426\%$ ) for the MEPS16 dataset using the SA algorithm.
+We observe the largest absolute improvement of $12.215\%$ in EOD ($20.641\%$ to $8.426\%$ ) for the MEPS16 dataset using SA.
 The largest relative improvement of $69.43\%$ ($8.962\%$ to $2.749\%$) was observed in the Default dataset using SA.
@@
 For Bank, the validation and test accuracies increase from 0.842 and 0.84 to 0.892 and 0.88 for SA and 0.882 and 0.881 for RW.
-The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is negative class and 1 is positive) as more and more neurons are dropped out from the DNN.
-The increase in negative predictions favors the true negatives as the datasets are higly imbalanced with more negative samples.
+The opposing trends of F1 score and accuracy can be attributed to the increase in negative predictions (0 is the negative class and 1 is positive) as neurons continue to be dropped out from the DNN.
+The increase in negative predictions favors the true negatives as the datasets are highly imbalanced with more negative samples.
 The Default, Bank, and MEPS16 datasets have 78%, 88%, and 83% of the data belonging to the negative class.
@@
 A high penalty multiplier would decrease the probability of state transitions, thereby encouraging the SA algorithm to exploit the search space near the current best state.
-The penalty multiplier ‘p’ directly affects the cost difference $\Delta E$ between states with acceptable and unacceptable F1 scores.
+The penalty multiplier p directly affects the cost difference $\Delta E$ between states with acceptable and unacceptable F1 scores.
 From Line 14 in Algorithm LABEL:alg:neuron_repair, we can see that a higher cost difference results in a lower transition probability.
-For a fixed temperature, a higher ‘p’ would increase the cost difference between good and bad states, thereby decreasing the probability of transitioning and encouraging exploitation.
-A lower ‘p’ would result in a lower cost difference and increase the transition probability, thereby encouraging exploration.
+For a fixed temperature, a higher p would increase the cost difference between good and bad states, thereby decreasing the probability of transitioning and encouraging exploitation.
+A lower p would result in a lower cost difference and increase the transition probability, thereby encouraging exploration.
 To determine the effect of the penalty multiplier, we run experiments with the Adult, MEPS16, Default, and Bank datasets with one seed by varying the F1 penalty multiplier $p\in\{0.5,1.0,1.5,2.0,2.5,3.0\}$ and F1 threshold multiplier to 0.98.
@@
 The time-out for SA runs is one hour.
-The brute force runs take 60 hours to explore the entire search space.
+The brute force exhaustive search takes 60 hours.
 Table 7 highlights the results of both search strategies.
@@
 The highest cost difference of $0.054$ shows the efficacy of randomized algorithms in identifying good states.
-SA is faster in reaching a good state compared to brute force as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost.
-Another reason for SA’s effectiveness is that it uses a probabilistic transition function that can transition to a state that has a worse cost and escape local minima.
-Gradient-based methods often struggle with the problem of getting stuck in local minima.
+SA is faster than brute force in reaching a good state as it follows a cost gradient to choose a path to a good state by always accepting states that improve the cost.
+Additionally, SA uses a probabilistic transition function that can transition to worse states and potentially escape local minima.
+Gradient-based methods often get stuck in local minima.
 Table 8 shows the distribution of the Best, Good, and Bad states.
 The best states are the states with the global optimal cost.
-The cost difference between the good and best states is less than 0.05.
+The cost of the good states is within 0.05 of the optimal cost.
 Bad states are states whose F1 scores are less than the 98% of the baseline F1 score.
@@
 This suggests that the good states occupy small “pockets” of the search space.
-The cost gradient leads SA to effectively identify such pockets.
-The probabilistic transition function enables SA to explore the region without getting stuck in local minima.
+The cost gradient effectively leads SA to identify such pockets and the probabilistic transitions prevent it from getting stuck in local minima.
 Limitations.
@@
 To address the internal validity and ensure our finding does not lead to an invalid conclusion, we follow the established SE guidelines and repeat the experiments 10 times, reporting both the average and 95% confidence intervals.
-In addition, we study the progress of algorithms during their runs, not just the final outcomes.
+In addition, we study the progress of algorithms during their runs, not just the outcomes.
 To ensure that our results are generalizable and address external validity, we perform our experiments on seven DNN benchmarks with various architectures.
@@
 Reweighting (Kamiran and Calders, 2012) augments the dataset where the data points in each group-label combination are weighted differently to ensure fairness.
-The reweighting process minimizes the discrepancy between the observed and expected probabilities of the favorable outcome occurring with the discriminated sensitive attribute.
+The reweighting process minimizes the discrepancy between the observed and expected probabilities of the favorable outcome occurring with the sensitive attribute.
 FairMask (Peng et al., 2022) uses the fact that protected attributes can be inferred by some combinations of other non-protected attributes (the “proxy” problem) to reduce the influence of protected attributes in the inference stage.
@@
 For the fairness model, they design a debugging strategy based on prior work (Wick et al., 2019; Chakraborty et al., 2021b) that attribute bias in training data to selection bias and label bias.
-LTDD (Li et al., 2022) hypothesizes that biased features in training data contributes to unfairness.
-To obtain unbiased features, they develop a linear regression based algorithm that measures the association between non-sensitive and sensitive features in the dataset.
+LTDD (Li et al., 2022) hypothesizes that biased features in training data contribute to unfairness.
+They develop a linear regression-based algorithm that measures the association between non-sensitive and sensitive features in the dataset to obtain unbiased features.
 The non-sensitive features are then modified to remove the biased parts to become independent of the sensitive features.
-After debiasing, the sensitive features are dropped from the dataset.
+The sensitive features are dropped from the training dataset after debiasing.
 The aforementioned works tamper with the precious training data samples that might lead to unrealistic representations of different protected groups.
@@
 B) In-processing techniques aim to improve fairness during training.
-One method, known as adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions.
-Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the learning objective (loss function), aiming to balance both accuracy and fairness.
-These strategies necessitate alterations in either the loss function or the parameters of the ML models.
+Adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions.
+Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the loss function to balance accuracy and fairness.
+These strategies necessitate alterations in either the loss function or the model parameters.
 Seldonian (Thomas et al., 2019; Metevier et al., 2019; Hoag et al., 2023) presented a technique that allows users to directly specify (arbitrary) undesirable behaviors as constraints and enforce them during training.
@@
 Parfait-ML (Tizpaz-Niari et al., 2022) is a gray-box evolutionary search algorithm that explores the ML hyperparameters to find configurations that minimize fairness while maintaining an acceptable accuracy.
-While mitigating unfairness during training might be less intrusive than changing the training datasets; they often require changing training algorithms that might be only effective for a particular algorithm and fairness definition.
-In addition, the training process is often expensive and might not be feasible for existing systems.
+While mitigating unfairness during training might be less intrusive than changing the training datasets, it often requires changing the training algorithms.
+This might only be effective for a particular algorithm and fairness definition.
+Additionally, the training process is often expensive and might not be feasible for existing systems.
 C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022).
-Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post-hoc correction mechanism that modifies the output of a biased binary classifier.
+Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post hoc correction mechanism that modifies the output of a biased binary classifier.
 EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor.
@@
 Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness.
-Care (Sun et al., 2022) presented Particle Swarm Optimisation (PSO) to repair DNNs.
+Care (Sun et al., 2022) presented a Particle Swarm Optimization (PSO) algorithm to repair DNNs.
 Faire (Li et al., 2023) proposes a method to repair unfair DNNs by altering the activations of neurons.
@@
 Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted.
-Compared to Faire, NeuFair does not retrain the model and does not separately use the protected attribute to identify unfair neurons.
+NeuFair, however, does not retrain the model and does not separately use the protected attribute to identify unfair neurons.
 We find an unfair subset of neurons given the entire feature vector, thereby accounting for the causal influence of the protected features on the non-protected features in the training dataset.
 Our analysis considers a group of neurons together rather than identifying the contributions of individual neurons.
-FairNeuron (Gao et al., 2022) proposes an algorithm to selectively retrain neurons in a neural network to mitigate unfairness.
+FairNeuron (Gao et al., 2022) proposes an algorithm to mitigate unfairness by selectively retraining neurons in a DNN.
 On a high level, the algorithm identifies input-output neuron paths corresponding to each sample in the training dataset that induces bias in the prediction.
-Then, they split the dataset into biased and unbiased subsets corresponding to the neuron activations in the paths.
-Finally, the algorithm re-trains the model with both training data splits by enabling random dropouts on the neurons in the biased paths and disabling any dropouts for neurons in the un-biased paths.
+Then, the dataset is split into biased and unbiased subsets corresponding to the neuron activations in the paths.
+Finally, the algorithm retrains the model with both training data splits by enabling random dropouts in the biased paths and disabling any dropouts in the unbiased paths.
 While FairNeuron utilizes dropouts to improve fairness, it fundamentally differs from NeuFair.
 First, we do not require access to the large training dataset, and we do not retrain the models.
-Second, NeuFair algorithms utilize deterministic dropout during inference time to mitigate unfairness with a flexible budget on the loss in F1 score, i.e., we identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly.
-Lastly, since we do not perform re-training, NeuFair does not update the weights of any neuron in the network.
+Second, NeuFair algorithms utilize deterministic dropout during inference time to identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly.
+Lastly, NeuFair does not retrain the DNN to update weights.
 8. Conclusion
-In this paper, we tackle the problem of mitigating unfairness in pre-trained deep neural networks using the dropout method.
+In this paper, we tackle the problem of mitigating unfairness in pre-trained DNNs using the dropout method.
 We showed that the neural dropout problem over the DNN models is computationally hard and presented NeuFair, a family of randomized algorithms to efficiently and effectively improve the fairness of DNNs.
@@
 For future work, there are a few exciting directions.
-First, instead of choosing the best state found during the search, we can analyze the top $k$ states found during the run.
+First, we can analyze the top $k$ states found during the run instead of choosing the best state.
 Such a strategy could reduce the difference in improvement observed between the validation and test sets.
@@
 Second, we plan to leverage dropouts in the fine-tuning process to understand the effects of randomized algorithms on mitigating the unfairness and toxicity of large language models.
-Lastly, alternate strategies like genetic algorithms can be considered to solve the combinatorial optimization problem of neuron dropout.
+Lastly, alternate strategies like genetic algorithms can be used to solve the combinatorial optimization problem of neuron dropout.
 We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems.
 The primary difference between SA/RW and genetic algorithms is that genetic algorithms create a population of candidate solutions at each step instead of a single solution.
-Genetic algorithms can be promising as they enable us to explore more states of the search space simultaneously.
-Data Availability. Our open-source tool NeuFair with all experimental subjects are publicly accessible: Link.
+This approach can be promising as it enables us to explore more states of the search space simultaneously.
+Data Availability
+Our open-source tool NeuFair with all experimental subjects are publicly accessible: https://github.com/vdasu/neufair.
+Acknowledgements.
+This material is based upon work supported by the National Science Foundation under Grant No. CNS-2230061.

### v3

@@
 Abstract.
-This paper investigates neuron dropout as a post-processing bias mitigation for deep neural networks (DNNs).
+This paper investigates neuron dropout as a post-processing bias mitigation method for deep neural networks (DNNs).
 Neural-driven software solutions are increasingly applied in socially critical domains with significant fairness implications.
-While neural networks are exceptionally good at finding statistical patterns from data, they may encode and amplify existing biases from the historical data.
+While DNNs are exceptional at learning statistical patterns from data, they may encode and amplify historical biases.
 Existing bias mitigation algorithms often require modifying the input dataset or the learning algorithms.
-We posit that the prevalent dropout methods that prevent over-fitting during training by randomly dropping neurons may be an effective and less intrusive approach to improve the fairness of pre-trained DNNs.
+We posit that prevalent dropout methods may be an effective and less intrusive approach to improve fairness of pre-trained DNNs during inference.
 However, finding the ideal set of neurons to drop is a combinatorial problem.
-We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference after training.
+We propose NeuFair, a family of post-processing randomized algorithms that mitigate unfairness in pre-trained DNNs via dropouts during inference.
 Our randomized search is guided by an objective to minimize discrimination while maintaining the model’s utility.
-We show that our design of randomized algorithms is effective and efficient in improving fairness (up to 69%) with minimal or no model performance degradation.
-We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of search algorithms on the results.
+We show that NeuFair is efficient and effective in improving fairness (up to 69%) with minimal or no model performance degradation.
+We provide intuitive explanations of these phenomena and carefully examine the influence of various hyperparameters of NeuFair on the results.
 Finally, we empirically and conceptually compare NeuFair to different state-of-the-art bias mitigators.
-conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austria
+Machine Learning, Bias Mitigation, AI Ethics
+copyright: acmlicenseddoi: 10.1145/3650212.3680380journalyear: 2024isbn: 979-8-4007-0612-7/24/09conference: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis; September 16–20, 2024; Vienna, Austriabooktitle: Proceedings of the 33rd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA ’24), September 16–20, 2024, Vienna, Austriasubmissionid: issta24main-p1278-pccs: Software and its engineering Search-based software engineeringccs: Computing methodologies Machine learning
 1. Introduction
@@
 Thus, the community presents various testing (Galhotra et al., 2017; Aggarwal et al., 2019; Zhang et al., 2020b), debugging (Gohar et al., 2023; Monjezi et al., 2023; Yu et al., 2023), and mitigation (Zhang and Harman, 2021; Biswas and Rajan, 2021) techniques to address fairness defects in data-driven software.
-Broadly, fairness mitigation can be applied in pre-processing stage (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing step (e.g., changing the loss function to include fairness constraints during training), or post-processing stage (e.g., changing the logic of a pre-trained model).
-However, when the decision logic of AI is encoded via deep neural networks, it becomes challenging to mitigate unfairness due to its black-box uninterpretable nature.
+Broadly, fairness mitigation can be applied in the pre-processing (e.g., increasing the representations of an under-represented group by generating more data samples for them), in-processing (e.g., changing the loss function to include fairness constraints during training), or post-processing (e.g., changing the logic of a pre-trained model) stage.
+However, when the decision logic of AI is encoded via DNNs, it becomes challenging to mitigate unfairness due to its black-box uninterpretable nature.
 We posit that a subset of neurons in a neural network disparately contributes to unfairness.
@@
 While the dropout strategy has been significantly used to prevent over-fitting during the training of DNNs, to the best of our knowledge, this is the first work to systematically leverage dropout methods as a post-processing bias mitigation method for improving the fairness of pre-trained DNN models.
-However, finding the optimal subset of dropout neurons is a combinatorial problem that requires an exhaustive search in the subset of all combinations of neuron dropouts, and hence is intractable.
+However, finding the optimal subset of dropout neurons is an intractable combinatorial problem that requires an exhaustive search over all possible combinations of neuron dropouts.
 To overcome the computational challenges, we explore the class of randomized algorithms with Markov chain Monte Carlo (MCMC) strategies to efficiently explore the search space with statistical guarantees.
 In doing so, we pose the following research questions:
-RQ1. How successful are randomized algorithms in fairness repairing of DNNs via dropouts?
+RQ1. How successful are randomized algorithms in repairing the unfairness of DNNs via dropouts?
 RQ2. Are there dropout strategies that improve model fairness and utility together?
@@
 Furthermore, we examine different hyperparameter configuration options of randomized algorithms.
-While some hyperparameters always influence fairness with positive or negative impacts; we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark.
+While some hyperparameters always influence fairness with positive or negative impacts, we detect a hyperparameter that defines a trade-off between explorations and exploitation that should be tuned as a constant variable in the cost function for each benchmark.
 Finally, we show the effectiveness of the SA algorithm, compared to the RW and the state-of-the-art post-processing (bias) mitigator (Monjezi et al., 2023).
@@
   3. (3)
-    Experimental evaluations. We implement the randomized algorithms in NeuFair and evaluate its effectiveness and efficiency vis-a-vis the state-of-the-art techniques.
+    Experimental evaluations. We implement the randomized algorithms in NeuFair and evaluate their effectiveness and efficiency vis-a-vis the state-of-the-art techniques.
 2. Background
-We provide a background on the model utility and fairness measures. Then, we briefly discuss various bias mitigation methods from the AI fairness literature.
+In this section, we provide a background on the various model utility and fairness metrics.
 2.1. Notions of Model Utility
@@
 Accuracy can be used to gauge the overall performance of a classifier.
-However, for imbalanced datasets used in fairness evaluations, accuracy is a poor metric as the number of negative samples far outweighs the positive samples.
-For example, predicting all samples as negative (0 true positives) in the Bank (Dua and Graff, 2017a) dataset yields an accuracy of 88% but the F1 score would be 0/undefined.
+However, accuracy is a poor metric for imbalanced datasets commonly used in fairness evaluations as the number of negative samples far outweighs the positive samples.
+For example, predicting all samples as negative (0 true positives) in the Bank (Dua and Graff, 2017a) dataset yields an accuracy of 88% but the F1 score would be 0 or undefined.
 Since F1 is defined as the harmonic mean of precision and recall, rather than the arithmetic mean, it penalizes performance significantly if either precision or recall is low.
@@
 Equal Opportunity is a relaxed variant of Equalized Odds with $Y=1$ (Hardt et al., 2016).
-In other words, Equalized Odds require the true positive and false positive rates to be equal across all sensitive groups.
-In contrast, Equal Opportunity only requires the true positive rate to be equal across all sensitive groups.
+Equalized Odds require the true positive and false positive rates to be equal across all sensitive groups.
+However, Equal Opportunity only requires the true positive rate to be equal across all sensitive groups.
 Demographic Parity is the weakest notion of fairness, and Equalized Odds is the strongest.
@@
 The DNN model is parameterized by the input dimension $m{+}r$, the output dimension $2$, the depth of hidden layers $n+1$, and the weights of its hidden layers $W_{0},W_{1},\ldots,W_{n}$.
-We describe the hidden layers with $\mathcal{M}\leftarrow[L_{I},L_{0},\ldots,L_{n},L_{O}]$, where $L_{I}$ and $L_{O}$ are the input and output layers, respectively, and ${L_{i}},\forall i\in[0,n]$, is the hidden layers.
+We describe the hidden layers with $\mathcal{M}\leftarrow[L_{I},L_{0},\ldots,L_{n},L_{O}]$, where $L_{I}$ and $L_{O}$ are the input and output layers, respectively, and ${L_{i}},\forall i\in[0,n]$, are the hidden layers.
 We assume that there exists a subset of neurons $\mathcal{N}\in{L_{i}},\forall i\in[0,n]$ in the hidden layers that disparately contribute to unfairness.
@@
 3.2. Inference Time Dropout for Fairness
-Dropout (Srivastava et al., 2014) is a technique proposed to improve the generalization of DNNs by preventing overfitting.
+Dropout (Srivastava et al., 2014) is a technique proposed to improve the performance of DNNs by preventing overfitting.
 Dropout sets all $w_{i}$ to 0 for a random set of neurons in the hidden layers with some probability during training.
 Once training is complete, dropout is not used, and all the neurons in the DNN are utilized to make predictions.
-While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the DNN during inference after training can significantly improve fairness with a minimal impact on model performance.
+While dropout has been traditionally used to prevent overfitting during training, we hypothesize that dropping neurons of the DNN during inference after training can significantly improve fairness with a minimal impact on performance.
 However, unlike traditional dropout, where a set of neurons are randomly dropped during training, we aim to identify a subset of neurons at the Pareto optimal curve of fairness-performance during inference.
-We consider a binary vector as the neuron state with $s=\{0,1\}^{N}$, where $N=\sum_{i=0}^{n}|L_{i}|$, $s_{i}$ indicates whether the neuron $i$ is dropped or not and $n$ is the number of layers.
-A pre-trained DNN model $\mathcal{D}$ does not include any dropouts, and hence all the indicators are 00.
+We consider a binary vector as the neuron state with $s=\{0,1\}^{N}$, where $N=\sum_{i=0}^{n}|L_{i}|$ and $s_{i}$ indicates whether the neuron $i$ is dropped or not and $n$ is the number of layers.
+A pre-trained DNN model $\mathcal{D}$ does not include any dropouts; hence, all the indicators are 00.
 Definition 3.0 (Desirable Dropout of Fairness vs. Utility).
@@
   1. RQ1
-    How successful are randomized algorithms in fairness repairing of DNNs via dropouts?
+    How successful are randomized algorithms in repairing the unfairness of DNNs via dropouts?
   2. RQ2
@@
 We evaluate NeuFair with five different datasets from fairness literature.
-For two of the datasets, we consider two different protected groups, which effectively results in a total of 7 different benchmarks.
-Table 5.3 presents of an overview of the datasets and protected groups.
+For two of the datasets, we consider two different protected groups, which effectively results in a total of 7 benchmarks.
+Table 5.3 presents an overview of the datasets and protected groups.
 The Adult Census Income (Dua and Graff, 2017b), Bank Marketing (Dua and Graff, 2017a), Compas Software (ProPublica, 2021), Default Credit  (Dua and Graff, 2017c), and Medical Expenditure (MEPS16) (for Healthcare Research and Quality, line) are binary classification tasks to predict whether an individual has income over $50$K, is likely to subscribe, has a low reoffending risk, is likely to default on the credit card payment, and is likely to utilize medical benefits, respectively.
@@
 For data preprocessing, we utilize standard techniques such as one-hot encoding for categorical features followed by min-max or standard scaling for numerical features.
-For the Compas dataset, we use a version of the dataset that has 12 features after feature selection that was used in (Tizpaz-Niari et al., 2022).
+For the Compas dataset, we use a version of the dataset used in (Tizpaz-Niari et al., 2022) that has 12 features after feature selection.
 The modified Compas dataset is available online [^4] https://github.com/Tizpaz/Parfait-ML/blob/main/subjects/datasets/compas.
@@
 5.2. Technical Details
-We run all our experiments on a desktop running Ubuntu 22.04.3 LTS with an Intel(R) Core(TM) i7-7700 CPU @ 3.60GHz processor, 32GB RAM, and a 1 TB HDD.
-The neural networks and repair algorithms were implemented using python3.8, torch==2.0.1,
+[…]
 numpy==1.24.3, and scikit-learn==1.3.1.
@@
 Unless otherwise specified, all runs reported in the paper have a time-out limit of $1$ hour with the following cost function:
-$C(s)=EOD_{s}+3.0\cdot EOD_{s_{0}}\cdot\mathbbm{1}(F1_{s}<(0.98\cdot F1_{s_{0}})).$
+$C(s)=EOD_{s}+3.0\cdot EOD_{s_{0}}\cdot\mathds{1}(F1_{s}<(0.98\cdot F1_{s_{0}})).$
 The F1 threshold is 98% of the validation F1 score of the unfair DNN, i.e., we tolerate a 2% degradation in the F1 score to improve fairness.
 A penalty of 3×3$\times$3 × the baseline unfairness is added to the fairness of the current state if its F1 score is less than the threshold.
-For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons depending on the overall size of the network.
+For all datasets, we set the minimum number of neurons to $n_{l}=2$ and vary the maximum number of neurons $n_{u}$ between $20\%-40\%$ of the number of hidden layer neurons.
 The $n_{u}$ values are 50, 24, 24, 20, and 135 for Adult, Compas, Bank, Default, and MEPS16 respectively.
@@
 Out of 70 runs (10 per benchmark), the SA and RW found a desirable state with 0.98×0.98$\times$0.98 × the initial F1 score in 69 cases.
-In addition, we re-run the experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$.
-The progress graphs of the RW strategy are “steep” initially and as the initial states are unfair and are heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold.
+In addition, we re-run the failed Default experiment with an F1 threshold multiplier of $0.96$ and identify a desirable state with a test set fairness of $6.26\%$, an improvement from $13.2\%$.
+The progress graphs of the RW strategy are “steep” initially as the initial states are unfair and are heavily penalized by the penalty multiplier as RW generates intermediate states with an F1 score less than the threshold.
 However, due to the cost function and probabilistic transitions, SA only explores worse states with a low probability.
-So, the progress graphs are not as steep initially.
+Thus, the progress graphs are not as steep initially.
 Table 3 and Table 4 highlight the mean and $95\%$ confidence intervals for the F1 score, accuracy, and EOD for all data subjects aggregated across 10 seeds.
@@
 Pathological Case.
-Table 5 highlights the EOD scores for COMPAS (Sex) with the original and repaired networks.
+Table 5 highlights the individual EOD scores for all seeds of COMPAS (Sex) with the original and repaired networks.
 Here, we observe that SA improves the EOD in 4 cases, does not improve in 1 case, and has a worse EOD in 5 cases.
@@
 First, there is a discrepancy in the original model statistics between the validation and test splits.
-The difference arises because the DNN tends to overfit slightly on the training/validation sets.
+The difference arises because the DNN tends to overfit slightly on the training and validation sets.
 This is evident from model metrics in Table 3.
@@
 During the search, the SA algorithm correctly identifies a desirable state according to the validation dataset and performs better than RW.
-However, given the low initial unfairness and the model’s tendency to perform better on the validation/training sets, we observe minor discrepancies in the test set.
+However, given the low initial unfairness and the model’s tendency to perform better on the training and validation sets, we observe minor discrepancies in the test set.
 The average unfairness, however, increases by 0.399% (2.522% to 2.921%), which is negligible.
@@
 The trained DNN models were optimized for F1 score on the validation dataset to improve performance on the underrepresented positive class.
-The accuracy increases as the increase in true negatives outweighs the drop in true positives.
-The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and false negatives increase.
+The accuracy increases as the increase in true negatives far outweighs the drop in true positives.
+The F1 score, however, always decreases as we lose precision and recall when the true positives decrease and the false negatives increase.
 Answer RQ2: The overall accuracy, as a model utility metric, may improve along with fairness using dropout strategies, as accuracy does not account for false positives and false negatives in imbalanced datasets.
@@
 We compare the efficacy of NeuFair against a state-of-the-art post-processing bias mitigator.
-Dice (Monjezi et al., 2023) used a fault localization technique via do logic to pinpoint a single neuron that significantly influences fairness.
+Dice (Monjezi et al., 2023) uses a fault localization technique via do logic to pinpoint a single neuron that significantly influences fairness.
 The approach computes the amounts of variation in the outcomes of each layer that depends on the sensitive attributes and de-activates a single neuron with the highest effects on fairness.
-Since the search space evaluates a single neuron, we brute-force the entire search space of one neuron dropout in a linear time.
+Since the search space evaluates a single neuron, we brute-force the entire search space of one neuron dropout in linear time.
 We note that the best improvement, feasible by Dice, cannot be more than the brute-force search.
@@
 We also note a similar trend with COMPAS (Sex), where the unfairness increases slightly.
-The largest discrepancy in fairness improvement between NeuFair and Dice is observed in the MEPS16 dataset, with an absolute difference of 10.778% and a relative improvement of 56.124% (19.204% and 8.426%) using the SA algorithm.
+The largest discrepancy in fairness improvement between NeuFair and Dice is observed in the MEPS16 dataset, with an absolute difference of 10.778% and a relative improvement of 56.124% (19.204% and 8.426%) using SA.
 Answer RQ4: NeuFair outperforms the state-of-the-art unfairness mitigation algorithm Dice (Monjezi et al., 2023) across all benchmarks. The relative improvement of NeuFair is 56% higher than Dice.
@@
 Bad states are states whose F1 scores are less than the 98% of the baseline F1 score.
-We observe that the best and good states are a negligible amount of the search space.
+We observe that the best and good states occupy a negligible amount of the search space.
 On the other hand, over 90% of the states are bad states.
@@
 Since we used randomized algorithms, our results might be sub-optimal.
-While simulated annealing provides statistical guarantees on the confidence and running time to reach the optimal state, in practice such guarantees might require multiple runs of algorithms over a long period of time.
+While SA provides statistical guarantees on the confidence and running time to reach an optimal state, in practice such guarantees might require multiple runs of algorithms over a long period of time.
 Additionally, we observe a discrepancy between the validation and test fairness improvements where the test set improves less than the validation set in both NeuFair algorithms.
@@
 The sensitive features are dropped from the training dataset after debiasing.
-The aforementioned works tamper with the precious training data samples that might lead to unrealistic representations of different protected groups.
+The works above tamper with the precious training data samples that might lead to unrealistic representations of different protected groups.
 Additionally, it may not always be feasible to modify training data.
-Our work is tangential as it focuses on post-processing mitigation that does not modify or access the training dataset and works on pre-trained unfair DNNs.
-B) In-processing techniques aim to improve fairness during training.
-Adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that not only improves prediction accuracy but also hinders the ability of adversaries to determine sensitive attributes from ML predictions.
+Our work is tangential, focusing on post-processing mitigation that does not modify or access the training dataset and works on pre-trained unfair DNNs.
+B) In-processing techniques improve fairness during training.
+Adversarial debiasing (Zhang et al., 2018), employs adversarial learning to develop a classifier that hinders the ability of adversaries to determine sensitive attributes from ML predictions.
 Another approach, the prejudice remover (Kamiran et al., 2012), incorporates a fairness-centric adjustment into the loss function to balance accuracy and fairness.
@@
 C) Post-processing techniques aim to modify the prediction outcomes of ML models to reduce discrimination (Hardt et al., 2016; Kamiran et al., 2012; Pleiss et al., 2017; Tao et al., 2022).
-Equalized Odds Processing (EOP) (Hardt et al., 2016) is a post hoc correction mechanism that modifies the output of a biased binary classifier.
+Equalized Odds Processing (EOP) (Hardt et al., 2016) modifies the output of a biased binary classifier to improve fairness.
 EOP solves an optimization problem using a linear program that utilizes the protected attribute, predictions, and true labels from a dataset to create an unbiased predictor.
-NeuFair, however, takes a fundamentally different approach to post-processing mitigation.
-First, NeuFair does not require training auxiliary models to mitigate unfairness.
-Second, NeuFair does not rely on randomization for fairness and generates a deterministic DNN classifier.
+NeuFair, however, does not require training auxiliary models to mitigate unfairness.
+Additionally, NeuFair does not rely on randomization for fairness.
 Predictions during inference time are deterministic after the neurons are dropped from the DNN.
-Lastly, Woodworth et al.
-(2017) shows that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data.
+Woodworth et al.
+(2017) also show that EOP requires the biased classifier to be Bayes optimal, which is practically impossible to learn from finite samples of data.
 Zhang and Sun (Zhang and Sun, 2022) adaptively intervened in the input data features and DNN internal neurons to improve fairness.
@@
 To create the two categories, Faire creates a clone of the unfair model and retrains it to predict the protected attribute.
-By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected.
+By comparing the activations of the corresponding neurons in both networks, Faire determines which neurons represent the protected and non-protected features.
 Finally, the protected neuron activations are penalized, and the non-protected neuron activations are promoted.
@@
 Finally, the algorithm retrains the model with both training data splits by enabling random dropouts in the biased paths and disabling any dropouts in the unbiased paths.
-While FairNeuron utilizes dropouts to improve fairness, it fundamentally differs from NeuFair.
-First, we do not require access to the large training dataset, and we do not retrain the models.
-Second, NeuFair algorithms utilize deterministic dropout during inference time to identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly.
-Lastly, NeuFair does not retrain the DNN to update weights.
+While FairNeuron uses dropout to improve fairness, it fundamentally differs from NeuFair.
+We do not retrain the DNN by using the training dataset.
+Also, NeuFair utilizes deterministic dropout during inference time to identify a desirable subset of neurons to drop that has a minimal impact on F1 score but the maximal impact on fairness, rather than dropping neurons randomly.
 8. Conclusion
@@
 Such a strategy could reduce the difference in improvement observed between the validation and test sets.
-We can extend to multi-valued protected attributes (age groups) or optimize for more than one protected attribute at a time.
-Second, we plan to leverage dropouts in the fine-tuning process to understand the effects of randomized algorithms on mitigating the unfairness and toxicity of large language models.
-Lastly, alternate strategies like genetic algorithms can be used to solve the combinatorial optimization problem of neuron dropout.
+We can extend to multi-valued protected attributes (e.g., age groups) or optimize for more than one protected attribute at a time (e.g., race and sex).
+Second, we plan to leverage dropouts in the fine-tuning process to understand the effect of randomized algorithms on mitigating the unfairness and toxicity of large language models.
+Lastly, alternate strategies like genetic algorithms may be used to solve the combinatorial optimization problem of neuron dropout.
 We use randomized algorithms for their simplicity in efficiently solving combinatorial optimization problems.
@@
 Data Availability
-Our open-source tool NeuFair with all experimental subjects are publicly accessible: https://github.com/vdasu/neufair.
+Our open-source tool NeuFair with all experimental subjects are publicly accessible on Zenodo (Dasu, 2024) and GitHub.
 Acknowledgements.
-This material is based upon work supported by the National Science Foundation under Grant No. CNS-2230061.
+This material is based upon work supported by the National Science Foundation under Grant No. CNS-2230060 and CNS-2230061.